what is the y-intercept of 3x+y=7

Found 2 solutions by terbear1963, checkley75:
Answer by terbear1963(43)

FIRST, YOU MUST WRITE THE EQUATION IN THE PROPER FORMAT. 3X+Y=7, MUST BE WRITTEN AS, Y=MX+B, WHERE MX IS THE SLOPE AND B IS THE Y INTERCEPT. SO, YOUR EQUATION MUST BE WRITTEN AS FOLLOWS : Y=-3X+7. YOUR Y INTERCEPT IS 7. I HOPE THIS HELPS YOU OUT SOME!? ANY OTHER MATH QUESTION, PLZ E MAIL ME AT cats7051@yahoo.com YOUR MATH FRIEND, TERRY

Answer by checkley75(3666)

3X+Y=7
Y=-3X+7
USING THE STANDARD LINE FORMULA OF:
Y=mX+b
WE GET THE SLOPE (m)=-3 & THE Y INTERCEPT (b=7).